“Ordering results for option prices”

Ludger Rüschendorf, University of Freiburg

In the first part of the talk we discuss option prices in incomplete models based on minimum distance martingale measures. For several classes of exponential semimartingale models one gets explicit formulas for the minimum distance martingale measures. There is a close relationship to utility based pricing. In the second part we establish some general criteria for ordering the option prices in exponential semimartingales by ordering conditions on the local characteristics of the model. Applications are discussed to diffusion models, stochastic volatility models, jump diffusions, and Lévy type exponential models.

/of*/Ru*/All*/Abst*/OrderingResults-041207.tex